OpenStudy (anonymous):

Write the standard equation of a circle that passes through (-5,5) with center (-10,-5).

OpenStudy (swamijatin2):

The (x - a) and (y - b) parts of the equation dictate the center of the circle, which is (a, b) = (-10, -5) (X - (-10)) = (x + 10) ( y - (-5)) = (y + 5) These sum of these 2 terms squared = r^2 Find the radius using Pythagorean formula: (-5 - (-10))^2 + (5 - (-5))^2 = 25 + 100 = 125 = r^2